What is the solubility of NH3 if it’s partial pressure is reduced to 220.0 mmgHg at 25 C


1
Expert's answer
2021-11-08T08:27:51-0500

"s
=\frac{
m
N
H
3}{
V}"

"\frac{51.8
 
g}{
100
 
m
l}
=
\frac{m
N
H
3}{
2
,
500}
 
m"


"m
N
H
3
=
\frac{51.8}{
 
g
100}
 
m
l
×
2
,
500
 
m
l"


mNH3=1,295 g